OpenStudy (anonymous):

Solve the following quadratic equation by using the quadratic formula 4x^2-3x-7=0

OpenStudy (er.mohd.amir):

ax^2+bx+c=0 x={-b+sqrt(b*b-4ac)}/2a x={-b-sqrt(b*b-4ac)}/2a here a=4,b=-3,c=-7